OpenStudy (anonymous):

Perpendicular to y=x+3

OpenStudy (anonymous):

Perpendicular values for lines are the inverse of y.

OpenStudy (anonymous):

"y = mx + c" where m is the gradient and c is the y intercept. Your equation: y = x + 3 A parallel line has the same gradient as the line it is parallel to. You need to know the points it passes through. Equation of a line is y -y1 = m(x -x1)

OpenStudy (anonymous):

the slope of a perpendicular line is the negative reciprocal of the slope of the original equation. If you are given a point, you can be able to complete the equation of the perpendicular line :)